# A Framework to Audit Email Address Privacy and Analyze Email Marketing Practices of Online Services and Apps

### v1

@@ 4.3. Temporal Analysis
The most relevant frequency for our analysis is the weekly cycle ($\approx 7$ days), denoted by $f=0.1436$ with a magnitude of $|X(f)|=393.39$. This weekly periodicity corresponds closely with the expected operational and behavioral cycles observed in email-sending patterns, reflecting regular weekly activity. We use the 7-day period as the seasonality figure for the time series decomposition. The decomposition highlights a linear, decreasing trend in the number of emails received over time. Additionally, it shows that the weekly seasonal component explains 1.17% of the variance in the number of emails received. The amount of seasonality in email sending behavior is mediated at the company level and will be explored in greater detail in following sections on the cross-sector and company-level analysis.
@@
Second, given that the observed weekly seasonality in sending patterns we look to specify the general time of day and day of week that email messages are sent. The polar heat maps shown in Figure 4 show that the most popular times messages are sent occur during the mid-week, Tuesday through Thursday, and emails are predominantly sent around midnight or during the afternoon work hours between 1300 and 1500 hours.
@@ 4.4. Clustering
These features were combined into a comprehensive feature set, standardized using StandardScaler, and subsequently reduced to five principal components using PCA. The dimensionality reduction to five principal components retained over 80% of the variance in the data. The PCA results show the factor loadings and below is an interpretation of the PCA component loadings:
@@
  • Component 3 reflects customer engagement through CRM and alert emails sent throughout the day and evening. Companies scoring high on this component focus on personalized communication and customer relationship management, engaging customers with CRM and alert messages at various times.
  • Component 4 emphasizes afternoon and evening email sending patterns. High scores on this component suggest companies that target specific times of the day, such as afternoon and evening hours, to optimize their communication strategies.
  • Component 5 highlights alert emails sent during non-standard hours, such as early morning and late evening. Companies with high scores on this component may send time-sensitive communications, such as urgent updates or notifications.
K-Means clustering was employed to group companies based on their PCA-transformed features. To determine the optimal number of clusters, we evaluated cluster validity using the Silhouette Score, a metric that measures how similar a point is to its assigned cluster compared to other clusters. We tested a range of cluster numbers (2-10) and identified an clustering configuration based on Silhouette Score and prior domain knowledge. For this analysis, a final model with five clusters was selected, as it effectively balanced intra-cluster coherence with inter-cluster separation. Using the PCA loadings, we interpret the characteristics of each cluster to better understand the distinct email marketing behaviors exhibited by companies within each group. The average scores of each cluster on the principal components highlight the dominant features that distinguish one cluster from another.
  • Cluster 0: Low-Activity Generalists
  • Cluster 1: Ultra High-Volume Promotional Marketers
  • Cluster 2: High-Volume Morning Promotional Marketers
    Cluster 2 is characterized by strong positive scores on Components 1 and 2 but negative scores on Components 3 and 4. This profile points toward companies that not only engage in frequent email activity but also strategically send emails during morning hours. The negative scores on Components 3 and 4 suggest less emphasis on CRM communications and afternoon or evening sends. These companies may leverage professional marketing services extensively to optimize their outreach.
  • Cluster 3: Diversified Marketers with Emphasis on CRM and Alerts
    Cluster 3 shows positive scores across all principal components, with particularly high scores on Components 2 and 3. These companies are distinguished by a strong focus on diversified email strategies, including CRM and alert-type communications sent throughout the day. The positive scores on Components 2 and 3 suggest these companies engage in targeted messaging at various times, aligning with customer-specific triggers or personalized engagement strategies. This cluster represents sophisticated marketers who prioritize customer engagement alongside promotional activities.
  • Cluster 4: Balanced Marketers with Strategic Timing
    Cluster 4 presents a moderate positive score on Component 1 and slight positive scores on Components 3, 4, and 5. This cluster captures companies that maintain a balanced approach to email marketing, with strategic timing and a mix of promotional and CRM-focused content. They are characterized by neither extremely high volume nor extreme specialization in any one type of content, suggesting a versatile and adaptive strategy. This cluster represents firms that aim for a well-rounded communication style, balancing frequency, content, and timing to optimize customer engagement without overwhelming recipients.
Figure 5. Cluster membership by top two principal components.
The clustering analysis, visually depicted in Figure 5, reveals distinct segments of companies based on their email marketing strategies, ranging from high-volume promotional tactics to CRM-focused approaches. The PCA loadings provide a clear view of the factors driving these segmentations.
  • Volume vs. Personalization Trade-off: Clusters 1 and 3 illustrate the trade-off between volume and personalization, with Cluster 1 companies prioritizing frequent promotional outreach, while Cluster 3 focuses on targeted CRM communication and customer engagement.
  • Professional Marketing Services Usage: Clusters 2 and 4 emphasize the strategic use of professional marketing tools, leveraging advanced platforms to optimize timing and content, indicative of a more sophisticated approach to email marketing.
  • Strategic Timing: Clusters 2, 3, and 4 highlight companies that not only consider what to send but also when to send, using timing as a strategic lever in their marketing mix.
@@ 4.5. Cross-Sector Differences
To explore potential differences in email sending behaviors across various sectors, we analyzed key metrics such as the total volume of emails sent, the types of emails (promotional, CRM, alert), and the timing of email sending (hourly and daily patterns). The Health sector was excluded from the analysis due to having only one company. Descriptive statistics for each sector were calculated to understand the average email sending behaviors (Table 1).
Table 1. Email volume and type by sector (averages).
Sector Total Promo CRM Alert
Consumer Discretionary 58.14 48.52 9.06 0.37
Financials 53.50 0.50 40.00 13.00
Communication Services 19.63 3.37 14.84 1.26
Information Technology 17.88 8.96 8.54 0.33
Email Volume and Type by Sector. A chi-squared test of independence was conducted to examine the relationship between email type and sector (Table 2). The test revealed a statistically significant association between email type and sector ($\chi^{2}(6)=1411.29,p<0.0001$), indicating that the distribution of email types varies by sector.
Table 2. Email types per sector.
Sector CRM Alert Promotional
Consumer Discretionary 571 23 3057
Communication Services 282 24 64
Information Technology 205 8 215
Financials 80 26 1
@@
  • The Consumer Discretionary sector, which includes companies like wish.com, wayfair.com, and kohls.com, predominantly sends promotional emails (83.73% of their emails). This aligns with strategies to stimulate immediate consumer purchases through promotions.
  • The Financials sector focuses heavily on CRM (74.77%) and alert emails (24.30%), reflecting the importance of customer relationship management and timely notifications in financial services.
  • The Communication Services sector sends a higher proportion of CRM emails (76.22%), suggesting an emphasis on maintaining customer relationships and providing service updates.
  • The Information Technology sector exhibits a more balanced distribution between promotional (50.23%) and CRMemails (47.90%), indicating a dual focus on marketing products and supporting existing customers.
Variability Among Individual Companies. A one-wayANOVA was conducted to compare the average emails sent per company by sector. The results indicated no statistically significant difference ($F(3,103)=1.4487,p=0.2330$), suggesting that the average email volume per company is similar across sectors. Similarly, the Kruskal-Wallis H tests for both hourly and daily email sending patterns revealed no statistically significant differences between sectors. This indicates that sectors do not differ significantly in terms of when they send emails.
The lack of significant differences in email volume and timing between sectors implies that we cannot generalize about an individual company’s email marketing sophistication or patterns based solely on its sector. Individual companies within the same sector may adopt diverse strategies in terms of email frequency and timing, reflecting company-specific marketing approaches.
However, the significant differences in email types suggest that sectors have characteristic focuses in their email content, even if the overall volume and timing are similar. This means that while the sector provides some insight into the nature of email communications (e.g., promotional vs. CRM focus), it does not predict how intensively or when individual companies will engage in email marketing.
Moreover, although we did not find broad evidence of sector-wide differences in email volume or timing, it is noteworthy that the most prolific email senders are from the Consumer Discretionary sector, particularly retail and product-oriented businesses. Companies like Wish, Wayfair, and Kohls are among the highest volume emailers, emphasizing promotional content to drive sales.
The statistical analysis indicates a significant association between email type and sector, highlighting that different industries utilize email communications to achieve their specific business objectives:
  • Retailers in the Consumer Discretionary sector engage in high-volume promotional emailing, aiming to drive sales through frequent promotions.
  • The Financials sector prioritizes CRM and alert emails, essential for customer engagement and timely financial updates.
  • The Communication Services sector focuses on CRM communications, possibly to maintain service satisfaction and customer loyalty.
  • The Information Technology sector maintains a balanced approach, combining promotional efforts with customer support communications.
These insights suggest that while the overall volume and timing of emails are consistent across sectors, the content and purpose of emails are tailored to industry-specific strategies.
@@ 5. Discussion
This study provides a comprehensive analysis of email marketing strategies employed by the most popular online websites and mobile applications. By collecting emails where each unique email address corresponds to a single company, we have ensured a precise lineage and isolation of source company to end user. This methodological approach, combined with a year-long email data collection period encompassing 111 unique companies, offers both breadth and depth to our dataset, distinguishing our work from prior studies.
Our research contributes to the ongoing conversation about how companies utilize digital channels to market to and maintain relationships with their customers. The extensive dataset allows us to observe real-world practices over an extended period, providing insights into not only the frequency and content of communications but also the strategies underlying them.
@@
The findings reveal no evidence of spam emails from third-party domains and a significant portion of consumer concerns regarding excessive promotional emails can be mitigated by opting out of marketing communications where possible. The data also shows that the bulk of unsolicited first-party emails originate from a small subset of companies, supporting the Pareto principle where approximately 80% of the emails come from 20% of the senders. This concentration suggests that aggressive email marketing is not a universal practice but is driven by specific companies, often within certain sectors like retail and consumer services.
@@
ASN Mapping Shows Reliance on Third Parties. A novel aspect of our study is the analysis of email provenance via the originating Autonomous System Numbers (ASNs). By tracing the sender IP addresses to their corresponding ASNs, we uncover the reliance of companies on third-party marketing and Customer Relationship Management (CRM) services. Major intermediaries such as Salesforce, Mailgun, SendGrid, and SparkPost are identified as key players facilitating the bulk of email marketing traffic.
@@
This approach sheds light on the ecosystem of email marketing infrastructure, highlighting how companies leverage specialized external platforms to execute their communication strategies. A complete picture, however, is elusive as many of the emails originate from public cloud providers such as AWS and GCP, so what third party is operating its service hosted on these hyperscalers isn’t able to easily be determined just from the IP address. It also raises questions about data privacy and the extent to which customer information is shared with third-party providers, a topic that warrants further investigation.
@@
The “What” and “When” Matter. Our temporal analysis indicates that while the overall volume and timing of emails are consistent across sectors, the content and purpose are tailored to industry-specific strategies. For instance, the Consumer Discretionary sector predominantly sends promotional emails, aligning with their goal of stimulating immediate purchases. In contrast, sectors like Financials focus on CRM and alert emails, reflecting the importance of customer engagement and timely information dissemination.
@@ 5.1. Legal and Ethical Considerations
The findings raise important legal and ethical considerations regarding compliance with privacy policies and marketing preferences. Despite regulations like the CAN-SPAM Act in the United States and the General Data Protection Regulation (GDPR) in the European Union, consumers continue to report receiving unsolicited emails. Our study suggests that opting out during registration can reduce the volume of such emails, but also highlights that some companies may not fully adhere to stated preferences.
@@
Policymakers and regulatory bodies may find these insights valuable in assessing the effectiveness of current regulations and identifying areas where enforcement or additional guidelines are necessary. Our study identifies that a vast majority of marketing emails are facilitated by third-party services such as Salesforce, Mailgun, and SendGrid, with 89.35% of emails being sent through the top 8 Autonomous System Numbers (ASNs). This centralization indicates potential focal points for regulatory oversight. Additionally, despite existing regulations like the CAN-SPAM Act and GDPR, our findings suggest that some companies may not fully comply with stated privacy policies and marketing preferences, as evidenced by the persistent high volume of unsolicited promotional emails from certain sectors. Policymakers should consider these patterns when evaluating regulatory frameworks, ensuring that opt-out mechanisms are straightforward, transparent, and honored in practice. Enhancing regulations to specifically reference and include the use of third-party marketing services and enforcing compliance can further improve consumer experiences and uphold consumer rights.
@@ 5.2. Limitations and Future Work
The richness of our dataset extends beyond metadata to include the full subject lines and body content of emails. This presents opportunities for future research into the specific language, persuasive techniques, and thematic elements employed in email marketing. To do this, future work can explore the content of emails in greater detail through natural language understanding (NLU) techniques. Additionally, most of the emails in the dataset contain images that could be rendered and then analyzed, yielding insights into specific persuasion techniques that are employed. Moreover, our dataset can support studies on the effectiveness of different email strategies by correlating communication patterns with customer engagement metrics, where available. Finally, investigating the role of cultural, regional, or demographic factors in email marketing practices also presents an avenue for further study. Future work could also investigate whether smaller companies, which may face less regulatory scrutiny compared to popular online services and apps, are more likely to share or leak users’ email addresses without consent.
A key component of our framework is its low cost of ongoing maintenance and scalability. Once deployed, it can monitor any number of online services and apps passively. A key advantage of this is that we can detect any future changes in the inbound email messages. So for example, if a company in the monitored set has a breach that leaks customer email addresses and then those addresses are used for spam campaigns or sold to unauthorized third parties, our system would be one of the first authoritative detection sources of such a breach.
@@ 6. Conclusion
This study offers a detailed examination of email address privacy and marketing practices among leading online services and apps, highlighting key patterns in communication strategies and the use of third-party services. Our findings contribute to a better understanding of how companies balance marketing objectives with consumer preferences and the implications for trust and engagement in the digital marketplace. By shedding light on these practices, we aim to inform both industry and policy discussions, promoting approaches that respect consumer rights while enabling effective marketing communications.
@@ Appendix A LLM Classification Workflow
Appendix B List of Domains
LABEL:table:all-domains shows the details and distribution of emails received from 111 different root domains.
Table 3. Distribution of emails received from 111 different root domains.
@@ Appendix B List of Domains
adobe.com Information Technology 0 37 24 13 0
airtable.com Information Technology 0 2 0 2 0
aliexpress.com Consumer Discretionary 0 37 33 3 1
apartmentfinder.com Consumer Discretionary 0 1 0 1 0
apartments.com Consumer Discretionary 0 8 1 7 0
autozone.com Consumer Discretionary 0 13 7 5 1
bandcamp.com Consumer Discretionary 0 4 1 3 0
bestbuy.com Consumer Discretionary 2 391 376 12 3

### v2

@@ 4.3. Temporal Analysis
The most relevant frequency for our analysis is the weekly cycle ($\approx 7$ days), denoted by $f=0.1436$ with a magnitude of $|X(f)|=393.39$. This weekly periodicity corresponds closely with the expected operational and behavioral cycles observed in email-sending patterns, reflecting regular weekly activity. We use the 7-day period as the seasonality figure for the time series decomposition.
The decomposition highlights a linear, decreasing trend in the number of emails received over time. Additionally, it shows that the weekly seasonal component explains only 1.17% of the variance in the number of emails received, which is evidence of varied strategies being employed in when emails are sent.
The amount of seasonality in email sending behavior is mediated at the company level and will be explored in greater detail in following sections on the cross-sector and company-level analysis. Given that the observed weekly seasonality in sending patterns we look to specify the general time of day and day of week that email messages are sent. The polar heat maps shown in Figure 6 show that the most popular times messages are sent occur during the mid-week, Tuesday through Thursday, and emails are predominantly sent around midnight or during the afternoon work hours between 1300 and 1500 hours, specifically targeting when users are most active online.
@@ 4.4. Clustering
These features were combined into a comprehensive feature set, standardized using StandardScaler, and subsequently reduced to five principal components using PCA, retaining over 80% of the variance in the data.
K-Means clustering was employed to group companies based on their PCA-transformed features. To determine the optimal number of clusters, we evaluated cluster validity using the Silhouette Score, a metric that measures how similar a point is to its assigned cluster compared to other clusters. We tested a range of cluster numbers (2-10) and identified that separating companies into two clusters yielded the highest silhouette score (0.831). Using the PCA loadings, we interpret the characteristics of each cluster to better understand the distinct email marketing behaviors exhibited by companies within each group. The average scores of each cluster on the principal components highlight the dominant features that distinguish one cluster from another.
  • Cluster 0: Low-Activity Generalists
  • Cluster 1: High-Volume Promotional Marketers
Figure 7. Cluster membership by top two principal components.
The clustering analysis, visually depicted in Figure 7, reveals distinct segments of companies based on their email marketing strategies, ranging from high-volume promotional tactics to CRM-focused approaches. The PCA loadings provide a clear view of the factors driving these segmentations. From this analysis, there are two clear factor loadings that drive this clustering:
@@
We see from this clustering that the dominant discriminator between these companies and their email sending patterns is the volume of specifically promotional emails that they are sending.
@@ 4.5. Cross-Sector Differences
To explore potential differences in email sending behaviors across various sectors, we analyzed key metrics such as the total volume of emails sent, the types of emails (promotional, CRM, alert), and the timing of email sending (hourly and daily patterns). Since companies are moving to integrate their sales and marketing channels (Trenz et al., 2020), and there are no uniformly accepted definitions that separate many of these companies (Lehrer and Trenz, 2022), we rely on broad categorizations without making definitive claims as to their generalizability. This analysis is to further the argument, supported by our data, that not all companies send these unsolicited marketing emails in the same manner, and that certain sectors may be greater offenders than others.
Email Volume and Type by Sector. A chi-squared test of independence was conducted to examine the relationship between email type and sector. The test revealed a statistically significant association between email type and sector ($\chi^{2}(14)=2138.858,p<0.0001$), indicating that the distribution of email types varies by sector, with results shown in Table 1.
Table 1. Sector email volume and type averages.
Sector Total Promo CRM Alert $n$
Omnichannel Retailers 139.56 133.22 5.89 0.44 9
E-tailer 123.23 104.69 17.31 0.38 13
Financial Services 53.50 0.50 40.00 13.00 2
Digital Services 23.69 7.89 15.17 0.60 35
Online Marketplaces 26.75 20.06 6.63 0.00 16
Communication Platforms 19.63 3.37 14.84 1.26 19
Online Entertainment 15.20 5.10 9.10 1.00 10
Brick & Mortar 19.00 14.60 4.00 0.40 5
@@
  • Retailers, both Omnichannel Retailers and online only E-tailers, predominantly send promotional emails. Examples of omnichannel companies include Kohls and BestBuy, while e-tailers include Wish.com and Wayfair.com.
  • The Financials sector, while only 2 of the 21 companies actually sent emails, those two focus heavily on CRM (75.4%) and alert emails (24.3%) more than any other sector. This could reflect those companies using emails for non-marketing activities like account notifications.
Variability Among Individual Companies. A one-wayANOVA was conducted to compare the average emails sent per company by sector. The results indicated statistically significant difference ($F(7,101)=3.5095,p=0.002$), suggesting that the average email volume per company differs across sectors. Kruskal-Wallis H tests for both hourly and daily email sending patterns revealed no statistically significant differences between sectors. This indicates that sectors do not differ significantly in terms of when they send emails.
The significant differences in email volume and types between sectors implies that we may be able to direct future work towards those companies and sectors that are specifically sending promotional advertisements to consumers–retail and product-oriented businesses. While individual companies within the same sector may adopt diverse strategies in terms of email frequency and timing, there are some clear patterns that different industries utilize email communications to achieve their specific business objectives.
@@ 5.1. New Insights
This study provides a comprehensive analysis of email marketing strategies employed by the most popular online websites and mobile applications. By collecting emails where each unique email address corresponds to a single company, we have ensured a precise lineage and isolation of source company to end user. This methodological approach, combined with a year-long email data collection period encompassing 150 apps and services, offers both breadth and depth to our dataset, distinguishing our work from prior studies. Our research contributes to the ongoing conversation about how companies utilize digital channels to market to and maintain relationships with their customers. The extensive dataset allows us to observe real-world practices over an extended period, providing insights into not only the frequency and content of communications but also the strategies underlying them.
@@
The findings reveal no evidence of UUSS emails from unknown or unauthorized third-parties and a significant portion of consumer concerns regarding excessive promotional emails can be mitigated by opting out of marketing communications where possible. The data also shows that the bulk of SOS emails originate from a small subset of companies, supporting the Pareto principle where approximately 80% of the emails come from 20% of the senders. This concentration suggests that aggressive email marketing is not a universal practice but is driven by specific companies, often within certain sectors like retail.
@@
ASN Mapping Shows Reliance on Authorized Third-Parties. A novel aspect of our study is the analysis of email provenance via the originating ASNs. By tracing the sender IP addresses to their corresponding ASNs, we uncover the reliance of companies on authorized third-party marketing and CRM services. Major intermediaries such as Salesforce, Mailgun, SendGrid, and SparkPost are identified as key players facilitating the bulk of email marketing traffic.
@@
This approach sheds light on the ecosystem of email marketing infrastructure, highlighting how companies leverage specialized external platforms to execute their communication strategies. A complete picture, however, is elusive as many of the emails originate from public cloud providers such as AWS and GCP, so what third-party is operating its service hosted on these hyperscalers isn’t able to easily be determined just from the IP address. It also raises questions about data privacy and the extent to which customer information is shared with third-party providers, a topic that warrants further investigation.
@@
The “What” and “When” Matter. Our temporal analysis indicates that while the overall volume and timing of emails are relatively consistent across sectors, the content and purpose are tailored to industry-specific strategies. For instance, retailers predominantly sends promotional emails, aligning with their goal of stimulating immediate purchases. In contrast, financial service companies and communication platforms focus on CRM and alert emails, reflecting the importance of customer engagement and timely information dissemination.
@@ 5.2. Legal and Ethical Considerations
The findings raise important legal and ethical considerations regarding compliance with privacy policies and marketing preferences. Despite regulations like the CCPA, CAN-SPAM and GDPR, consumers continue to report receiving unsolicited emails. Our study suggests that opting out during registration can reduce the volume of such emails, but also highlights that some companies may not fully adhere to opt-out preferences.
@@
Policymakers and regulatory bodies may find these insights valuable in assessing the effectiveness of current regulations and identifying areas where enforcement or additional guidelines are necessary. Our study identifies that a vast majority of marketing emails are facilitated by third-party services such as Salesforce, Mailgun, and SendGrid, with 89.35% of emails being sent through the top 8 ASNs. This centralization indicates a potential focal point for improving regulatory oversight. Policymakers should consider these patterns when evaluating regulatory frameworks, ensuring that opt-out mechanisms are straightforward, transparent, and honored in practice. Enhancing regulations to specifically reference and include the use of third-party marketing services and enforcing compliance can further improve consumer experiences and uphold consumer rights.
@@ 5.3. Limitations and Future Work
The richness of our dataset extends beyond metadata to include the full subject lines and body content of emails. This presents opportunities for future research into the specific language, persuasive techniques, and thematic elements employed in email marketing. To do this, future work can explore the content of emails in greater detail through Natural Language Understanding (NLU) techniques (Allen, 1988; Canonico et al., 2018). Additionally, most of the emails in the dataset contain images that could be rendered and then analyzed (Zhang et al., 2023; Fei et al., 2024), yielding insights into specific persuasion techniques that are employed. Moreover, our dataset can support studies on the effectiveness of different email strategies by correlating communication patterns with customer engagement metrics, where available. Finally, investigating the role of cultural, regional, or demographic factors in email marketing practices also presents an avenue for further study. Future work could also investigate whether smaller companies, which may face less regulatory scrutiny compared to popular online services and apps, are more likely to share or leak users’ email addresses without consent.
While our study analyzes the correlation between spam reports and IP usage, identifying potential IP hopping behaviors, it does not yet evaluate whether such practices align with the companies’ stated privacy policies. Future work can focus on bridging this gap by cross-referencing email-sending behaviors with documented privacy commitments. This analysis would help determine if companies engaging in practices like IP hopping are acting inconsistently with their stated policies, providing deeper insights into accountability and compliance within the email ecosystem.
A key component of our framework is its low cost of ongoing maintenance and scalability. Once deployed, it can monitor any number of online services and apps passively. A key advantage of this is that we can detect any future changes in the inbound email messages. So for example, if a company in the monitored set has a breach that leaks customer email addresses and then those addresses are used for spam campaigns or sold to unauthorized third-parties, our system would be one of the first authoritative detection sources of such a breach.
@@ 6. Conclusion
This study offers a detailed examination of email inbox privacy and marketing practices among leading online services and apps, highlighting key patterns in communication strategies and the use of third-party services. Our findings contribute to a better understanding of how companies balance marketing objectives with consumer preferences and the implications for trust and engagement in the digital marketplace. By shedding light on these practices, we aim to inform both industry and policy discussions, promoting approaches that respect consumer rights while enabling effective email marketing communications.
@@ Appendix A LLM Classification Workflow
Appendix B List of Received Email Domains
LABEL:table:all-domains shows the details and distribution of emails received from 109 different root domains.
Table 2. Distribution of emails received from 109 different root domains.
@@ Appendix B List of Received Email Domains
adobe.com Digital Services 0 37 24 13 0
airtable.com Digital Services 0 2 0 2 0
aliexpress.com E-tailer 0 37 33 3 1
apartmentfinder.com Online Marketplace 0 1 0 1 0
apartments.com Online Marketplace 0 8 1 7 0
autozone.com Brick and Mortar 0 13 7 5 1
bandcamp.com Digital Services 0 4 1 3 0
bestbuy.com Omnichannel 1 391 376 12 3